40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Thu Aug 12 18:32:37 2010, Alastair, Laser, GYRO, Beam overlap etc TEK00000.PNGTEK00001.PNGTEK00002.PNG
    Reply  Fri Aug 13 15:50:39 2010, Jenna, Laser, GYRO, Beam overlap etc 
Message ID: 927     Entry time: Thu Aug 12 18:32:37 2010     Reply to this: 932
Author: Alastair 
Type: Laser 
Category: GYRO 
Subject: Beam overlap etc 

We started out today by trying to improve the beam overlap.  After talking to Aidan, and with Koji and Rana's suggestions, we began by beating the beams on the BS at the output.  We replaced the lens before the PD with a ~3cm focal length one.  Then we used the adjuster on the BS itself to make the beams overlap at the PD.  We got the signal to increase up to -31dBV (this was about 2% contrast).

We tried to improve this by adjusting the overlap by altering the BS, and one of the steering mirrors before the BS, but we couldn't improve it much.

Based on the 20mV signal from the DC input it seemed like we were only getting 27uW out of our 50uW on the actual photodiode.  It is difficult to get the lens exactly the correct distance from the PD because we have to focus down so small (it's impossible to see how big the beam is using an IR card or viewer).  We tried mounting the lens on a translation stage and maximising the power on the PD.  We managed to get about 60% increase in power doing this.

Looking at the 95MHz signal on the spectrum analyzer the frequency shifts back and forward quite a bit (~100kHz).  This goes away if you turn off the frequency modulation input on the VCO driving the AOM.  Looking at the CW beam (the AOM direction beam) on the monitor the mode appears to be a little unstable even though the CCW beam is not.  At this point we started looking at the reflection locking path for the AOM.  The PD for this locking scheme was getting a bit saturated so we installed an 80% BS to reduce the power on the PD down to about 1.5mW.

At this point we looked at the errror signal from the CW direction while sweeping the laser.  It looked a bit strange (see TEK00000.PNG  below.  The blue trace is the PD signal and the yellow is the error signal) and after making sure the beam was on the PD correctly we decided to check the polarization of the beam going into the cavity.  We found that it had become slightly rotated.  It probably got altered when we were aligning the beam through the Faraday.  We rotated it back to S and checked that the  cavity was still locking.  We put a power meter on the transmission side and maximized the power transmitted by steering the beam into the cavity.  The max we got was CCW 2.21mW out from 7.4mW in.  We checked the CW polarization which was still correct, and also maximized it's power out (this side does not have a good mode matching solution in place) and we are getting 4.5mW our of 25.9mW.  Since the output beams were now a little uneven in power we rotated the polarizer that controls the relatvie power in the two arms such that we are now getting ~12uW of power on the PD from each direction.

The error signal from the CW side looked much improved after changing the polarization (TEK00001.PNG) but the CW mode still looked unstable.  When we lock both directions of the cavity and look at these signals again we see this (TEK00002.PNG) oscillation in the signal.  It is massive.  It basically takes up the full error signal and is no doubt the reason why the mode looks unstable.

If you turn off the FM input on the VCO driving the AOM then the oscillation is 50% smalller but still there.  We wondered if this was coming from the Tektronix, so we swapped the AOM over to the Marconi, but we see exactly the same noise with this if the FM deviation is set to ~300kHz.  It does go lower if we lower the deviation setting on the Marconi below about 100kHz.  This might be the noise we're seeing in the 95MHz signal in transmission.

**Edit** we just spoke to Frank who thinks this may be because we have the gain too high in this loop (though turning down the PDH box to almost zero didn't get rid of the oscillation).  Tomorrow we'll try reducing this further.

Attachment 1: TEK00000.PNG  17 kB  | Hide | Hide all
TEK00000.PNG
Attachment 2: TEK00001.PNG  18 kB  | Hide | Hide all
TEK00001.PNG
Attachment 3: TEK00002.PNG  40 kB  Uploaded Thu Aug 12 19:32:52 2010  | Hide | Hide all
TEK00002.PNG
ELOG V3.1.3-